If the demand function faced by a firm is:

Q = 90 – 2P

TC = 2 + 57Q – 8Q2 + Q3


Determine the level of output at which the firm maximizes the profit.



1
Expert's answer
2021-04-30T11:00:39-0400

i) . The firm’s profit-maximizing output level is:

MR = MC,

MR = TR'(Q) = 45 - Q,


"MR=\frac {\Delta TR} {\Delta Q}"


"TR=Price \u00d7Quantity (P\u00d7Q)"


"MC=TC \n\u2032\n (Q)=3Q \n^2\n \u221216Q+57"


"3Q \n^2\n \u221216Q+57=45\u2212Q"


"3Q \n^2\n \u221215Q+12=0"


"Q^2 - 5Q + 4 = 0"


"Q_1 = 4 units, Q_2 = 1 unit."